OpenStudy (anonymous):

8. You are using 0.05 gallons per mile of a SUV gas tank that holds 16 gallons. Make a T-table for the equation for this problem, using the x-values 0, 160, 320. Type in the ordered pairs you get as your answer. 9. You are using 0.05 gallons per mile of a SUV gas tank that holds 16 gallons. Use the graph you made for the equation for this problem, to approximate the answer to the following question: You drove 300 miles after filling up. How much gas was left in the tank, according to your graph?

OpenStudy (anonymous):

8. y=16-0.05x (plug values 0,160, and 320 in for x) y=16-0.05(0)=16, y=16-0.05(160)=8, y=16-0.05(320)=0 therefore pairs are (0,16), (160,8), (320,0) 9. y=16-0.05(300) = 1 ...therefore 1 gallon left.

OpenStudy (anonymous):

Will help me with three more questions please,

OpenStudy (anonymous):

ya post them